△ABC lies in the structural support system of the Ferris wheel.
If m∠A = 30° and AB = AC = 25 ft,
find the measures (in degrees) of ∠B and ∠C.
BASE ANGLE (Angle B) + VERTEX ANGLE (ANGLE A)+ SECOND BASE ANGLE (Angle C) = 180 and the fact that we again have an isosceles triangle, making Angle B and Angle C the same measure.

Answer by ikleyn(52787)

.

In this problem,


    A + B + C = 180  degrees,

    B = C

    A = 30 degrees.


Therefore

   30 + 2B = 180

        2B = 180 - 30

        2B =    150

         B =    150/2 = 75.


ANSWER.  B = C = 75 degrees.   Angles B and C are congruent and each of them has the measure of 75 degrees.

Solved.